Net Present Value Method and Present Value Index

MVP Sports Equipment Company is considering an investment in one of two machines. The sewing machine will increase productivity from sewing 160 baseballs per hour to sewing 290 per hour. The contribution margin is $0.44 per baseball. Assume that any increased production of baseballs can be sold. The second machine is an automatic packing machine for the golf ball line. The packing machine will reduce packing labor cost. The labor cost saved is equivalent to $21 per hour. The sewing machine will cost $325,300, have a seven-year life, and will operate for 1,400 hours per year. The packing machine will cost $100,100, have a seven-year life, and will operate for 1,200 hours per year. MVP seeks a minimum rate of return of 12% on its investments.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine the net present value for the two machines. Use the table of present values of an annuity of $1 above. Round to the nearest dollar.

Sewing Machine Packing Machine
Present value of annual net cash flows $ $
Less amount to be invested $ $
Net present value $ $

b. Determine the present value index for the two machines. If required, round your answers to two decimal places.

Sewing Machine Packing Machine
Present value index

c. If MVP has sufficient funds for only one of the machines and qualitative factors are equal between the two machines, in which machine should it invest? (If both present value indexes are the same, either machine will grade as correct.)
SelectPacking MachineSewing MachineItem 9
Answer #1
Annual net cash flows Sewing Machine =(290-160)*0.44*1400= $80080
Annual net cash flows Packing Machine =1200*21= $25200
a
Sewing Machine Packing Machine
Present value of annual net cash flows 365485 115013
Less amount to be invested 325300 100100
Net present value 40185 14913
b
Present value index:
Sewing Machine 1.12 =364485/325300
Packing Machine 1.15 =115013/100100
c
Packing Machine, as its Present value index is higher